# Dynamical tide modified Roche limit in eccentric, asynchronous binaries

### v1

@@ 4 Conclusion and discussion
In this work, we studied the nonlinear hydrodynamical stability of a donor $M$ in a compact orbit around a point-particle companion $M^{\prime}$. Our study considers both the exact nonlinear equations derived from the affine model (P. Diener et al., 1995), which self-consistently apply to homogeneous, incompressible ellipsoids, and the expanded equations following T. Van Hoolst (1994), which can be readily extended to realistic stars and planets. Our study extends the classic Roche limit derived by S. Chandrasekhar (1963) of a donor in hydrostatic equilibrium with circular orbit and synchronous rotation to the fully dynamical regime, allowing for arbitrary orbital eccentricities and donor rotation rates (though restricted to aligned spins). Schematically, the finite-frequency corrections to the equilibrium tide (accurate for low eccentricities; §2.2) can be captured by the toy model in Eq. (5). The dynamical tide excited in more eccentric systems (§2.3) corresponds to the toy model of Eq. (6). More specifically, the Roche limit for moderately eccentric systems ($e\lesssim 0.3$) is described by our Fig. 2. While a small asynchronicity tends to decrease $D_{p}^{(\rm th)}/R_{t}$ (i.e., making the donor more tightly bound) by reducing the centrifugal force, eccentricity can dominate and make $D_{p}^{(\rm th)}/R_{t}$ greater (donor more subject to mass loss) than the prediction of S. Chandrasekhar (1963) by the finite-frequency corrections. For a more eccentric binary with $e\sim 0.9$, the fluid perturbation can grow chaotically due to nonlinear interactions at the pericenter. This eventually leads to a runaway in mode amplitude at a pericenter separation $\gtrsim 30\%$ greater than the Roche limit of circular orbits with synchronous rotations (Figs. 6 vs. 1). Our theoretical result explains the numerical simulation of J. Guillochon et al. (2011) on the high-eccentricity migration of gaseous planets, and allows the result to be physically extrapolated across possible parameter space and to other similar systems, e.g., rpTDEs. This study also provides a new mechanism to trigger and maintain the chaotic tide that is of lower order in $(1/D_{p})$ than all previously considered mechanisms (§3.1).
@@
We mainly focused on a homogeneous, incompressible ellipsoid to assist a direct comparison with S. Chandrasekhar (1963). This simple model is also sufficient to illustrate the crucial role of dynamical effects in the Roche problem and its corrections to the commonly used fit of P. P. Eggleton (1983). We established that predictions based on the exact nonlinear equations of an affine ellipsoid can be largely reproduced by the expanded equations to the three-wave order, or $\mathcal{O}(\xi^{3})$ in $\Delta E_{\rm mode}$. Including the next order improves the accuracy, but does not modify the qualitative features (Fig. 1). While the affine model of P. Diener et al. (1995) allows for different values of $n$ and $\Gamma$, the starting point, Eq. (12), is not self-consistently satisfied by a general fluid. Nonetheless, the expanded equations of T. Van Hoolst (1994) will apply generically; see also N. N. Weinberg et al. (2012); N. N. Weinberg (2016). This allows the theoretical study of the Roche limit to be extended to realistic astrophysical objects. It also allows other modes, such as g-modes (present if $\Gamma>1+1/n$), or those with $l\geq 3$ (important for binaries with $M^{\prime}\sim 1$) to be included in the problem. Based on the affine model, an increasing $n$ reduces the single-passage $D_{p}^{(\rm th)}/R_{t}$ (P. Diener et al., 1995). It would be interesting to revisit this prediction quantitatively with generic shapes of $\boldsymbol{\xi}$ (non-affine) under (T. Van Hoolst, 1994) and consider the multi-orbit effects.
@@
In fact, the set of equations required to derive the Roche limit has been used in H. Yu et al. (2023) to study the frequency shift of the f-mode due to the companion’s tidal field in coalescing binary neutron stars (see also T. Pitre & E. Poisson 2025 for its relativistic extension). The two effects share the same physical origin. A detailed comparison with H. Yu et al. (2023) will enable us to quantify the errors of applying the affine model to $n>0$ fluids.
@@
The analysis of H. Yu et al. (2023) also provides further support for why we can truncate the analysis at the three-wave order for the Roche limit. Specifically, in this work, the interacting $m=2$ and $m=0$ modes are both directly linearly driven by the tide, allowing them to directly interact at the three-wave order. This is of the same order as the frequency shift of H. Yu et al. (2023) and lower order than the anharmonicity and pg-instability (N. N. Weinberg et al., 2013; N. N. Weinberg, 2016; T. Venumadhav et al., 2014), which fundamentally originate from four-wave interaction ($\mathcal{O}(\xi^{3})$ in the equation of motion). In the pg-instability, the p-mode is not linearly forced by the tide, but instead nonlinearly driven through the three-wave channel by an f-mode and a g-mode, schematically written as $q_{p}\simeq\kappa q_{f}q_{g}$ with the subscript indicating the mode (see appendix A of N. N. Weinberg et al. 2013). The p-mode then back-reacts on the g-mode, as
@@
causing the instability when $\omega_{g,{\rm eff}}^{2}<0$ or $\kappa q_{f}q_{f}^{\ast}>1$. Because the three-wave interaction is used twice with the p-mode acting only as a mediator, the pg-three-wave interaction comes in the same order as a direct four-wave interaction between two f-modes and two g-modes, corresponding to a forcing term of the form $\eta q_{f}^{\ast}q_{f}q_{g}$ in the equation above. See also similar discussions about the nonlinear anharmonicity studied in H. Yu et al. (2021) and K. J. Kwon et al. (2024, 2025) caused by effective four-wave interactions. In this work, however, the three-wave interaction is only used once because the $m=0$ mode has a linearly forced component ($q_{0}\sim q_{2}$ here instead of $q_{0}\sim\epsilon q_{2}$ as in the pg-instability), making it a lower-order effect than the anharmonicity. This can also be seen from the left panel of fig. B1 in H. Yu et al. (2023); compare specifically the gray (equivalent to the instability of this work) and the yellow (equivalent to the pg-instability/anharmonicity) curves. This theoretical argument is reflected in Fig. 1, and is numerically justified by Figs. 3 and 5. Including the four-wave interaction (Eqs. A33a and A33b) modifies the threshold $D_{p}^{\rm(th)}/R_{t}$ by only a few percent (Fig. 1).
@@
Similar to S. Chandrasekhar (1963), our study does not directly address the evolution of a fluid element after a mode’s amplitude runs away. Two crucial questions remain to be addressed. The first is how much mass is removed from $M$ and at what rate. Whereas in circular or low-eccentricity cases, a steady stream of mass flow from the donor to the accretor may form (F. H. Shu & S. H. Lubow, 1981), in the highly eccentric cases, the mass loss should occur as discrete bursts whenever the wave breaks (after the mode amplitude runs away). An upper bound on the mass loss $\Delta M$ can be set based on the energy released, $\Delta E_{\rm mode}\sim 0.1$. If most of the energy is converted to heat (M. MacLeod et al., 2022), then $\Delta M$ from the wave breaking may be much smaller than $\Delta E_{\rm mode}$. Nonetheless, a large amount of heat may inflate the donor, making it more likely to lose mass in the subsequent evolution, which quickly disrupts the donor (H. Yu & F. Dai, 2024). On the other hand, if the donor’s density increases as it loses mass, it may survive the mass loss (S.-F. Liu et al., 2013). One such example could be planets in the HN desert (D. J. Armstrong et al., 2020). The second question is how a mass element lost from the donor affects the orbit. The numerical simulation of J. Guillochon et al. (2011) suggests that the total mass and angular momentum of the system are likely to be non-conserved when the mass loss happens with a highly eccentric orbit. One may study the effects on the orbital dynamics following, e.g., S.-S. Huang (1963); J. F. Sepinsky et al. (2009); A. S. Hamers & F. Dosopoulou (2019); S. Y. Lau & H. Yu (2025). Alternatively, since the exterior potential of a perturbed, oscillatory star is simply $\propto Y_{2m}/D^{3}$, it is straightforward to find a test particle’s trajectory under a Roche potential modified by the quadrupolar potential of the donor (and potentially the accretor). The trajectory will determine the fate of the ejected mass and its impact on the orbit. Of particular interest are the changes in the binary’s total mass and orbital angular momentum. We plan to study these in follow-up investigations.
@@
We thank Fei Dai and Zhen Pan for useful discussions during the preparation of this manuscript. This work is supported by Montana NASA EPSCoR Research Infrastructure Development under award No. 80NSSC22M0042 and NSF grant No. PHY-2308415.
@@ Appendix A Equations of motion of the affine model
where $\boldsymbol{x}$ and $\boldsymbol{r}$ are the original and perturbed locations of a mass element. The Lagrangian displacement is denoted by $\boldsymbol{\xi}$, which is further decomposed into eigenmodes with eigenfunctions $\boldsymbol{\xi}_{m}$ according to Eqs. (7) and (8). As argued in the main text, inertial modes and $m=\pm 1$ f-modes are both ignored in this work. Because $\xi_{m}$ is linear in the original position, perturbations of a homogeneous ellipsoid are therefore affine as in B. Carter & J. P. Luminet (1983, 1985); P. Diener et al. (1995). We additionally introduce a trace term $\boldsymbol{\xi}_{\rm tr}=q_{\rm tr}\boldsymbol{x}$, corresponding to a monopolar (radial with $l=0$) perturbation. While a homogeneous ellipsoid has no radial modes in the linear problem, we find $q_{\rm tr}$ is needed to account for rotation and nonlinear interactions starting at the four-wave order. Note that the trace ($l=0$) mode is normalized differently from the $l=2$ ones; $q_{\rm tr}=1$ corresponds to a surface displacement of unity. This allows us to rewrite
@@
The equation of motion is obtained from variations of the energies. While the affine model is self-consistently satisfied only for an incompressible, homogeneous fluid with a polytropic index $n=0$ and adiabatic index $\Gamma=\infty$, we nonetheless follow (P. Diener et al., 1995) (see also B. Carter & J. P. Luminet 1983, 1985) and allow the energies to be described by a generic pair of $(n,\Gamma)$, without requiring buoyancy neutral (i.e., $\Gamma$ can differ from $(1+1/n)$), while still assuming the perturbed fluid is described by Eq. (A2). The error due to this inconsistency will be assessed in future studies through direct comparisons with T. Van Hoolst (1994).
@@
where $\mathcal{M}_{\ast}=\int x_{i}x_{i}dM/3$. Numerically, $\mathcal{M}_{\ast}=1/5$ for $n=0$ and $\mathcal{M}_{\ast}=0.1306910$ for $n=1$. Using the first law of thermodynamics $\Delta(dU/dM)=(P/\rho^{2})\Delta\rho$ and $\rho(r)=(\det[\boldsymbol{q}])^{-1}\rho(x)$, the variation of the internal energy is
$\Delta U$ $=\frac{1-\det[\boldsymbol{q}]}{\det[\boldsymbol{q}]}\Pi,$ (A8)
$\text{where }\Pi$ $=\int\frac{P}{\rho}dM=\frac{(\det[\boldsymbol{q}])^{1-\Gamma}}{5-n}.$ (A9)
The explicit form of $\Delta U$ is readily obtained by plugging in $\det[\boldsymbol{q}]$ from Eq. (A3), and its derivative with respect to $q_{ij}$ is $\partial U/\partial q_{ij}=-\Pi(\boldsymbol{q}^{-1})_{ji}$. This will be used to derive the equation of motion in terms of $q_{ij}$. To assist comparisons with T. Van Hoolst (1994), we also provide a series expansion in terms of $q_{m}$’s, as
$\Delta U$ $\simeq-\frac{3}{5-n}q_{\rm tr}+\frac{25q_{0}^{2}+50q_{2}q_{-2}+24(3\Gamma-1)q_{\rm tr}^{2}}{8(5-n)}$
$\quad-\frac{125\sqrt{6}q_{0}^{3}+450(6\Gamma-1)q_{0}^{2}q_{\rm tr}-750\sqrt{6}q_{0}q_{2}q_{-2}+900(6\Gamma-1)q_{2}q_{-2}q_{\rm tr}+72(9\Gamma(3\Gamma-1)+2)q_{\rm tr}^{3}}{144(5-n)}+$... (A10)
We have verified that the series matches exactly eq. (19) of T. Van Hoolst (1994) for $n=0$, in which case $\boldsymbol{\xi}_{m}$ are given by Eq. (8).
@@
From the definition, it is easy to see that the total gravitational binding energy $\Omega$ is given by the trace of $\Omega_{ij}$, $\Omega={\rm Tr}[\Omega_{ij}]$. Because the equations of motion depends on $\partial\Omega/\partial q_{ij}=-\Omega_{ik}(\boldsymbol{q}^{-1})_{jk}$, we compute the full $\Omega_{ij}$ matrix here. While B. Carter & J. P. Luminet (1983) gave a formula for $\Omega_{ij}$ in their eq. (6.15), it involves solving a complicated integral in their eq. (6.4) whose explicit solution is non-trivial to obtain. We instead follow S. Chandrasekhar & N. R. Lebovitz (1962) to obtain a closed-form solution. For this, it is convenient to first consider the potentials in the barred frame obtained by rotating the system azimuthally by an angle $\phi_{q}$ via the $\boldsymbol{R}$ matrix. In this system, the potential energy tensor is diagonal and given by
@@
where $\bar{I}_{ij}$ are still given by Eq. (A6), and we explicitly include the 5/(5n)5/(5-n) factor to generalize the result to arbitrary values of $n$ (with the caveat that for $n>0$, $\boldsymbol{r}$ is not strictly affine). The one-index symbols $A_{i}$ are defined in eq. (5) of S. Chandrasekhar & N. R. Lebovitz (1962) and computed explicitly in terms of $a_{i}$ in their eqs. (15)-(17). For completeness, we reproduce them below. Define
@@
Note that the results are obtained by swapping the subscripts $(2,3)$ in the corresponding equations in S. Chandrasekhar & N. R. Lebovitz (1962). This is because in our coordinate system, $a_{1}>a_{3}>a_{2}$ most of the time, though occasionally, it is possible to have $a_{2}>a_{3}$. In this case, Eq. (A14) still holds as long as one lets $\sin^{2}\theta_{e}>1$ and uses the reciprocal modulus transformation of elliptical integrals to handle $m>1$ (P. F. Byrd & M. Friedman, 1954)
@@
where we again expanded the exact solution in powers of $q_{m}$ in the last equality to assist with a direct comparison with T. Van Hoolst (1994). For a homogeneous ellipsoid, discontinuities in the derivatives of the gravitational potentials at the surface are non-trivial to evaluate, and we present an explicit evaluation of those terms in Appendix B.
@@
Note that we work in the frame corotating with the donor and use $\Phi_{C}=\Phi-\omega t$ to account for the Doppler effect. In comparison to, e.g., eq. (B1) of P. Diener et al. (1995), we additionally included the centrifugal force due to a rotation $\omega$ (assumed to be along the $z$ axis, also the direction of the orbital angular momentum), giving rise to an energy
@@
By further contracting Eq. (A22) with the symmetric-tracefree tensor $\mathcal{Y}_{lm}^{ij}$ (E. Poisson & C. M. Will, 2014), with
@@
Since the affine model is self-consistent only for homogeneous ellipsoids with $(n=0,\Gamma=\infty)$, we discuss the equations of motion for this case with further details. First note that the natural frequency of the $l=0$ mode is $\omega_{\rm tr}^{2}=(3\Gamma-4)/(5-n)$, which approaches infinity as $\Gamma\to\infty$. Therefore, we can approximate $\ddot{q}_{\rm tr}=0$, allowing us to solve for the trace part as
@@
We have expanded the results in $1/\Gamma$ and kept only the leading-order terms. Substituting them back into the equations of motion for quadrupolar modes, we obtain the complete description of the dynamics for $(n,\Gamma)=(0,\infty)$ at the four-wave order up to the ignorance of the Coriolis force (which can be included at the leading order following Eqs. 14-16; H. Yu & S. Y. Lau 2025) which reads

### v2

@@ IV Conclusion and discussion
In this work, we studied the nonlinear hydrodynamical stability of a donor $M$ in a compact orbit around a point-particle companion $M^{\prime}$. Our study considers both the exact nonlinear equations derived from the affine model (Diener et al., 1995), which self-consistently apply to homogeneous, incompressible ellipsoids, and the expanded equations following Van Hoolst (1994), which can be readily extended to realistic stars and planets. Our study extends the classic Roche limit derived by Chandrasekhar (1963) of a donor in hydrostatic equilibrium with circular orbit and synchronous rotation to the fully dynamical regime, allowing for arbitrary orbital eccentricities and donor rotation rates (though restricted to aligned spins). Schematically, the finite-frequency corrections to the equilibrium tide (accurate for low eccentricities; §II.2) can be captured by the toy model in Eq. (6). The dynamical tide excited in more eccentric systems (§II.3) corresponds to the toy model of Eq. (7). More specifically, the Roche limit for moderately eccentric systems ($e\lesssim 0.3$) is described by our Fig. 2. While a small asynchronicity tends to decrease $D_{p}^{(\rm th)}/R_{t}$ (i.e., making the donor more tightly bound) by reducing the centrifugal force, eccentricity can dominate and make $D_{p}^{(\rm th)}/R_{t}$ greater (donor more subject to mass loss) than the prediction of Chandrasekhar (1963) by the finite-frequency corrections. For a more eccentric binary with $e\sim 0.9$, the fluid perturbation can grow diffusively due to nonlinear interactions at the pericenter. This eventually leads to a runaway in mode amplitude at a pericenter separation $\gtrsim 30\%$ greater than the Roche limit of circular orbits with synchronous rotations (Figs. 6 vs. 1). Our theoretical result explains the numerical simulation of Guillochon et al. (2011) on the high-eccentricity migration of gaseous planets, and allows the result to be physically extrapolated across possible parameter space and to other similar systems, e.g., rpTDEs. This study also provides a new mechanism to trigger and maintain the diffusive (chaotic) tide that is of lower order in $(1/D_{p})$ than all previously considered mechanisms (§III.1).
@@
We mainly focused on a homogeneous, incompressible ellipsoid to assist a direct comparison with Chandrasekhar (1963). This simple model is also sufficient to illustrate the crucial role of dynamical effects in the Roche problem and its corrections to the commonly used fit of Eggleton (1983). We established that predictions based on the exact nonlinear equations of an affine ellipsoid can be largely reproduced by the expanded equations to the three-wave order, or $\mathcal{O}(\xi^{3})$ in $\Delta E_{\rm mode}$. Including the next order improves the accuracy, but does not modify the qualitative features (Fig. 1). While the affine model of Diener et al. (1995) allows for different values of $n$ and $\Gamma$, the starting point, Eq. (13), is not self-consistently satisfied by a general fluid. Nonetheless, the expanded equations of Van Hoolst (1994) will apply generically; see also Weinberg et al. (2012); Weinberg (2016). This allows the theoretical study of the Roche limit to be extended to realistic astrophysical objects. It also allows other modes, such as g-modes (present if $\Gamma>1+1/n$), or those with $l\geq 3$ (important for binaries with $M^{\prime}\sim 1$) to be included in the problem. Based on the affine model, an increasing $n$ reduces the single-passage $D_{p}^{(\rm th)}/R_{t}$ (Diener et al., 1995). It would be interesting to revisit this prediction quantitatively with generic shapes of $\boldsymbol{\xi}$ (non-affine) under (Van Hoolst, 1994) and consider the multi-orbit effects.
@@
In fact, the set of equations required to derive the Roche limit has been used in Yu et al. (2023) to study the frequency shift of the f-mode due to the companion’s tidal field in coalescing binary neutron stars (see also Pitre and Poisson 2025 for its relativistic extension). The two effects share the same physical origin. In particular, our analysis also predicts an effective frequency shift of the $m=2$ f-mode given by (cf. Eq. 21 but allowing for arbitrary $n$ using results from the Appendix A)
$\frac{\Delta(\omega_{f}^{2})}{\omega_{f}^{2}}=[\frac{95\sqrt{6}}{84}q_{0}-\frac{5}{8}(5-n)\mathcal{M}_{\ast}]\simeq\frac{65}{28}(5-n)\mathcal{M}_{\ast}\epsilon,$ (35)
where the linear, static limit of $q_{0}\simeq-\sqrt{6}(5-n)\mathcal{M}_{\ast}/4$ is used. This agrees with eq. 16 of Yu et al. (2023) in both the functional form and the numerical value, validating the affine approximation.
@@
The analysis of Yu et al. (2023) also provides further support for why we can truncate the analysis at the three-wave order for the Roche limit. Specifically, in this work, the interacting $m=2$ and $m=0$ modes are both directly linearly driven by the tide, allowing them to directly interact at the three-wave order. This is of the same order as the frequency shift of Yu et al. (2023) and lower order than the anharmonicity and pg-instability (Weinberg et al., 2013; Weinberg, 2016; Venumadhav et al., 2014), which fundamentally originate from four-wave interaction ($\mathcal{O}(\xi^{3})$ in the equation of motion). In the pg-instability, the p-mode is not linearly forced by the tide, but instead nonlinearly driven through the three-wave channel by an f-mode and a g-mode, schematically written as $q_{p}\simeq\kappa q_{f}q_{g}$ with the subscript indicating the mode (see appendix A of Weinberg et al. 2013). The p-mode then back-reacts on the g-mode, as
@@
causing the instability when $\omega_{g,{\rm eff}}^{2}<0$ or $\kappa q_{f}q_{f}^{\ast}>1$. Because the three-wave interaction is used twice with the p-mode acting only as a mediator, the pg-three-wave interaction comes in the same order as a direct four-wave interaction between two f-modes and two g-modes, corresponding to a forcing term of the form $\zeta q_{f}^{\ast}q_{f}q_{g}$ in the equation above. See also similar discussions about the nonlinear anharmonicity studied in Yu et al. (2021) and Kwon et al. (2024, 2025) caused by effective four-wave interactions. In this work, however, the three-wave interaction is only used once because the $m=0$ mode has a linearly forced component ($q_{0}\sim q_{2}$ here instead of $q_{0}\sim\epsilon q_{2}$ as in the pg-instability), making it a lower-order effect than the anharmonicity. This can also be seen from the left panel of fig. B1 in Yu et al. (2023); compare specifically the gray (equivalent to the instability of this work) and the yellow (equivalent to the pg-instability/anharmonicity) curves. This theoretical argument is reflected in Fig. 1, and is numerically justified by Figs. 3 and 5. Including the four-wave interaction (Eqs. A33a and A33b) modifies the threshold $D_{p}^{\rm(th)}/R_{t}$ by only a few percent (Fig. 1).
@@
Similar to Chandrasekhar (1963), our study does not directly address the evolution of a fluid element after a mode’s amplitude runs away. Two crucial questions remain to be addressed. The first is how much mass is removed from $M$ and at what rate. Whereas in circular or low-eccentricity cases, a steady stream of mass flow from the donor to the accretor may form (Shu and Lubow, 1981), in the highly eccentric cases, the mass loss should occur as discrete bursts whenever the wave breaks (after the mode amplitude runs away). An upper bound on the mass loss $\Delta M$ can be set based on the energy released, $\Delta E_{\rm mode}\sim 0.1$. If most of the energy is converted to heat (MacLeod et al., 2022), then $\Delta M$ from the wave breaking may be much smaller than $\Delta E_{\rm mode}$. Nonetheless, a large amount of heat may inflate the donor, making it more likely to lose mass in the subsequent evolution, which quickly disrupts the donor (Yu and Dai, 2024). On the other hand, if the donor’s density increases as it loses mass, it may survive the mass loss (Liu et al., 2013). One such example could be planets in the HN desert (Armstrong et al., 2020). The second question is how a mass element lost from the donor affects the orbit. The numerical simulation of Guillochon et al. (2011) suggests that the total mass and angular momentum of the system are likely to be non-conserved when the mass loss happens with a highly eccentric orbit. One may study the effects on the orbital dynamics following, e.g., Huang (1963); Sepinsky et al. (2009); Hamers and Dosopoulou (2019); Lau and Yu (2025). Alternatively, since the exterior potential of a perturbed, oscillatory star is simply $\propto Y_{2m}/D^{3}$, it is straightforward to find a test particle’s trajectory under a Roche potential modified by the quadrupolar potential of the donor (and potentially the accretor). The trajectory will determine the fate of the ejected mass and its impact on the orbit. Of particular interest are the changes in the binary’s total mass and orbital angular momentum. We plan to study these in follow-up investigations.
@@
We thank Eric Coughlin, Yanqin Wu, Brad Hansen, Smandar Naoz, Grant Weldon, Fei Dai and Zhen Pan for useful discussions during the preparation of this manuscript. This work is supported by Montana NASA EPSCoR Research Infrastructure Development under award No. 80NSSC22M0042 and NSF grant No. PHY-2308415.
@@ Appendix A Equations of motion of the affine model
where $\boldsymbol{x}$ and $\boldsymbol{r}$ are the original and perturbed locations of a mass element. The Lagrangian displacement is denoted by $\boldsymbol{\xi}$, which is further decomposed into eigenmodes with eigenfunctions $\boldsymbol{\xi}_{m}$ according to Eqs. (8) and (9). As argued in the main text, inertial modes and $m=\pm 1$ f-modes are both ignored in this work. Because $\xi_{m}$ is linear in the original position, perturbations of a homogeneous ellipsoid are therefore affine as in Carter and Luminet (1983, 1985); Diener et al. (1995). We additionally introduce a trace term $\boldsymbol{\xi}_{\rm tr}=q_{\rm tr}\boldsymbol{x}$, corresponding to a monopolar (radial with $l=0$) perturbation. While a homogeneous ellipsoid has no radial modes in the linear problem, we find $q_{\rm tr}$ is needed to account for rotation and nonlinear interactions starting at the four-wave order. Note that the trace ($l=0$) mode is normalized differently from the $l=2$ ones; $q_{\rm tr}=1$ corresponds to a surface displacement of unity. This allows us to rewrite
@@
The equation of motion is obtained from variations of the energies. While the affine model is self-consistently satisfied only for an incompressible, homogeneous fluid with a polytropic index $n=0$ and adiabatic index $\Gamma=\infty$, we nonetheless follow (Diener et al., 1995) (see also Carter and Luminet 1983, 1985) and allow the energies to be described by a generic pair of $(n,\Gamma)$, without requiring buoyancy neutral (i.e., $\Gamma$ can differ from $(1+1/n)$), while still assuming the perturbed fluid is described by Eq. (A2). The error due to this inconsistency will be assessed in future studies through direct comparisons with Van Hoolst (1994).
@@
where $\mathcal{M}_{\ast}=\int x_{i}x_{i}dM/3$. Numerically, $\mathcal{M}_{\ast}=1/5$ for $n=0$ and $\mathcal{M}_{\ast}=0.1306910$ for $n=1$.
The gas energy is $U=\int(\epsilon/\rho)dM$, where $\epsilon$ and $\rho$ are respectively the internal energy density and mass density. We assume that the gas follows a polytropic equation of state, where
$P=(\Gamma-1)\epsilon\propto\rho^{\Gamma},$ (A8)
with $P$ the pressure and $\Gamma=(\partial\log P/\partial\log\rho)_{\rm entropy}$ is the adiabatic index. Using further $\rho(\boldsymbol{r})=\rho(\boldsymbol{x})/\det[q]$, we have
$\Pi=\int\frac{P}{\rho}dM=\frac{(\det[q])^{1-\Gamma}}{5-n},\text{ and }U=\frac{\Pi}{\Gamma-1}.$ (A9)
The explicit form of $U$ is readily obtained by plugging in $\det[\boldsymbol{q}]$ from Eq. (A3), and its derivative with respect to $q_{ij}$ is $\partial U/\partial q_{ij}=-\Pi(\boldsymbol{q}^{-1})_{ji}$. This will be used to derive the equation of motion in terms of $q_{ij}$.
To assist comparisons with Van Hoolst (1994), we also provide a series expansion in terms of $q_{m}$’s, as
$U$ $\simeq\frac{1}{(5-n)(\Gamma-1)}-\frac{3}{5-n}q_{\rm tr}+\frac{25q_{0}^{2}+50q_{2}q_{-2}+12(3\Gamma-2)q_{\rm tr}^{2}}{8(5-n)}$
$\quad-\frac{125\sqrt{6}q_{0}^{3}-750\sqrt{6}q_{0}q_{2}q_{-2}+450(3\Gamma-1)q_{0}^{2}q_{\rm tr}+36(3\Gamma-1)[25q_{2}q_{-2}+2q_{\rm tr}^{2}(3\Gamma-2)]q_{\rm tr}}{144(5-n)}+$... (A10)
We have verified that the series matches exactly eq. (19) of Van Hoolst (1994) for $n=0$, in which case $\boldsymbol{\xi}_{m}$ are given by Eq. (9).
@@
From the definition, it is easy to see that the total gravitational binding energy $\Omega$ is given by the trace of $\Omega_{ij}$, $\Omega={\rm Tr}[\Omega_{ij}]$. Because the equations of motion depends on $\partial\Omega/\partial q_{ij}=-\Omega_{ik}(\boldsymbol{q}^{-1})_{jk}$, we compute the full $\Omega_{ij}$ matrix here. While Carter and Luminet (1983) gave a formula for $\Omega_{ij}$ in their eq. (6.15), it involves solving a complicated integral in their eq. (6.4) whose explicit solution is non-trivial to obtain. We instead follow Chandrasekhar and Lebovitz (1962) to obtain a closed-form solution. For this, it is convenient to first consider the potentials in the barred frame obtained by rotating the system azimuthally by an angle $\phi_{q}$ via the $\boldsymbol{R}$ matrix. In this system, the potential energy tensor is diagonal and given by
@@
where $\bar{I}_{ij}$ are still given by Eq. (A6), and we explicitly include the 5/(5n)5/(5-n) factor to generalize the result to arbitrary values of $n$ (with the caveat that for $n>0$, $\boldsymbol{r}$ is not strictly affine). The one-index symbols $A_{i}$ are defined in eq. (5) of Chandrasekhar and Lebovitz (1962) and computed explicitly in terms of $a_{i}$ in their eqs. (15)-(17). For completeness, we reproduce them below. Define
@@
Note that the results are obtained by swapping the subscripts $(2,3)$ in the corresponding equations in Chandrasekhar and Lebovitz (1962). This is because in our coordinate system, $a_{1}>a_{3}>a_{2}$ most of the time, though occasionally, it is possible to have $a_{2}>a_{3}$. In this case, Eq. (A14) still holds as long as one lets $\sin^{2}\theta_{e}>1$ and uses the reciprocal modulus transformation of elliptical integrals to handle $m>1$ (Byrd and Friedman, 1954)
@@
where we again expanded the exact solution in powers of $q_{m}$ in the last equality to assist with a direct comparison with Van Hoolst (1994). For a homogeneous ellipsoid, discontinuities in the derivatives of the gravitational potentials at the surface are non-trivial to evaluate, and we present an explicit evaluation of those terms in Appendix B.
@@
Note that we work in the frame corotating with the donor and use $\Phi_{C}=\Phi-\omega t$ to account for the Doppler effect. In comparison to, e.g., eq. (B1) of Diener et al. (1995), we additionally included the centrifugal force due to a rotation $\omega$ (assumed to be along the $z$ axis, also the direction of the orbital angular momentum), giving rise to an energy
@@
By further contracting Eq. (A22) with the symmetric-tracefree tensor $\mathcal{Y}_{lm}^{ij}{\color[rgb]{0,0,0}=(\mathcal{Y}_{l,-m}^{ij})^{\ast}}$ (Poisson and Will, 2014), with
@@
Since the affine model is self-consistent only for homogeneous ellipsoids with $(n=0,\Gamma=\infty)$, we discuss the equations of motion for this case with further details. First note that the natural frequency of the $l=0$ mode is $\omega_{\rm tr}^{2}=(3\Gamma-4)/(5-n){\color[rgb]{0,0,0}/\mathcal{M}_{\ast}}$, which approaches infinity as $\Gamma\to\infty$. Therefore, we can approximate $\ddot{q}_{\rm tr}=0$, allowing us to solve for the trace part as
@@
We have expanded the results in $1/\Gamma$ and kept only the leading-order terms. Substituting them back into the equations of motion for quadrupolar modes, we obtain the complete description of the dynamics for $(n,\Gamma)=(0,\infty)$ at the four-wave order up to the ignorance of the Coriolis force (which can be included at the leading order following Eqs. 15-17; Yu and Lau 2025) which reads